solve each system by grapting
1. 4y=x+2
2x+y=5





: solve each system by grapting
1. 4y=x+2
2x+y=5






Answer by checkley77(3848)

4y=x+2
y=x/4+2/4
y=x/4+.5 (red line)
2x+y=5
y=2y-5 (green line).
 graph( 300, 200, -6, 5, -10, 10, x/4 +.5, 2y -5) (graph 300x200 pixels, x from -6 to 5, y from -10 to 10, of TWO functions x/4 +.5 and 2y -5).